OpenStudy (anonymous):

A recipe that makes 7 servings calls for 1/2 cup of apple juice and 2/3 cup of cranberry juice. How many cups of juice are there per serving?

OpenStudy (igreen):

1/2 + 2/3 = 1 1/6. Divide by 7: 1 1/6 / 7 = 1/6. So there are 1/6 cups of juice per serving.

OpenStudy (anonymous):

thanks :)